OpenStudy (anonymous):

2m+n=-5 2m+3n=2 use elimination method by addition or subtraction

OpenStudy (anonymous):

change x to m and n to y 3 different ways here! Solution: x = 3.25, y = -1.5. Explanation: Solving: 2x+y = 5 2x+3y = 2 Using Substitution Method. Expressing x as a function of y (from the first equation) x = (5 -y)/2 Substituting this expression into the second equation: 2(5+y)/2+3y = 2 Simplifying this expression: 4y = -6 or y = -1.5 Now, substituting y = -1.5 back into the x expressed as a function of y: x = (5-y)/2 = 3.25